# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### 16 Dec 2023 v2

@@ 3.2 Relationships between variables
Figure 3: Regressions between four different variables and social health outcomes. From left to right, the x-axes show human likeness, consciousness, experience, and agency. The y-axis shows social health. Each point represents one participant. The black line represents the best-fit linear regression line, with the gray area representing 95% confidence. Graphs 3A-D show companion bot user data, and graphs E-H show the control group’s data. Graphs A and E show the relationship between the human-likeness index and social health index, graphs B and F show the relationship between the consciousness index and social health index, C and G show the relationship between the experience index and social health index, and D and H show the relationship between the agency index and social health index. For R, $R^{2}$, and statistical significance, see text.
@@
One possible explanation for why users have a more positive perception of companion bots than nonusers is as follows. People may be more disturbed or threatened by companion bots as they perceive them to be more conscious, agentic, experiential, and humanlike. Those who have more experience with companion bots may find them to be less humanlike and more artificial, more of a technological tool than a threat to human social roles, thereby reducing their concerns and increasing their perception of a social health benefit. To test this possible explanation, we examined the relationships between variables measured in this study. Figure 3 shows the relationships between five different indices. We tested the internal consistency for each of the five indices used in our analysis (social health, consciousness, experience, agency, and human likeness), and each index had good reliability (Cronbach’s alpha>0.8). Figure 3A, showing data from the companion bot users, illustrates the relationship between the social health index (based on an average of questions 3-5) and the human-likeness index (based on an average of questions 22-28). The graph shows a positive correlation (R=0.52, b=0.45, $R^{2}$=0.26) that was statistically significant in a regression analysis (F=21.93, P<0.0001). Companion bot users who thought that the chatbot was more humanlike also found it to be more beneficial to social health, in direct contradiction to the hypothesis.
@@
Figure 3B shows a similar positive and significant correlation between the social health index and the consciousness index (based on an average of questions 12-16, assessing how much the subjects thought the chatbot had a conscious mind) (R=0.44, b=0.23, $R^{2}$=0.18, F=16.32, P<0.001). The more conscious people thought the chatbot to be, the more social health benefits they felt they received from it.
@@
Figure 3C shows a positive and significant correlation between the social health index and the experience index (based on an average of questions 6-11, assessing how much the subjects thought the chatbot had a subjective experience) (R=0.43, b=0.24, $R^{2}$=0.17, F=15.33, P<0.001). The more subjective experience people thought the chatbot to have, the more social health benefits they felt they received from it.
@@
Similarly, Figure 3D shows a positive and significant correlation between the social health index and the agency index (based on an average of questions 4-10, assessing how much the subjects thought the chatbot had active agency) (R=0.32, b=0.24, $R^{2}$=0.09, F=7.869, P<0.01). The more agentic people thought the chatbot to be, the more social health benefits they felt they received from it.
@@
Previous researchers suggested that perceptions of consciousness, experience, and agency can be considered overlapping but still partially distinct subcategories that help make up the larger construct of human likeness [38-40]. Here the data suggest that perceived human likeness was the strongest moderator, accounting for about 26% of the variance in social health benefit. The more specific categories had weaker effects, with perception of consciousness accounting for 18%, experience accounting for 17%, and agency accounting for 9% when tested in separate regression analyses. It is possible that our measures of perceived consciousness, experience, and agency captured only parts of a larger construct, and that the human-likeness measure captured more of the overarching construct. It may also be that it is simply easier for subjects to rate human likeness than to rate subtler, hidden properties like perceived consciousness.
@@
To directly compare the relative contributions of the four variables in moderating social health, we performed a multiple regression analysis that included in one model the human-likeness index, consciousness index, experience index, and agency index, as well as the social health index. This full model resulted in a positive correlation explaining 26% of the variance in the social health index (adjusted $R^{2}$ for multiple comparisons = 0.26). We then removed variables one at a time, to determine which variable, when removed, caused the greatest drop in model performance. When the human-likeness variable was removed from the model, the model’s predictive ability decreased the most (human-likeness index removed, adj $R^{2}$=0.17; consciousness index removed, adj $R^{2}$=0.26; experience index removed, adj $R^{2}$=0.27; agency index removed, adj $R^{2}$=0.27). The data indicate that mind perception moderates social health benefits, with perception of human likeness being the most predictive variable.
@@
Figure 3E, F, G, and H show the corresponding results for the control group of people who did not use a companion bot. Remarkably, the relationships are nearly the same as the user group’s. For example, Figure 3E shows how subjects rated the likely social health benefits of the companion bot as a function of their judgement of how humanlike the companion bot was. On average, social health benefits were reported as lower by the control group (Figure 3E) compared to the companion bot user group (Figure 3A). Control subjects had a generally more pessimistic view of the companion bot. Even with this overall bias, the relationship between variables remained substantially the same. As control subjects judged the bot to be more humanlike, they also judged that it would have a greater benefit on their social health (R=0.50, b=0.48, $R^{2}$=0.24, F=38.96, P<0.0001). Smaller but still significant relationships were found between the social health index and the consciousness index, experience index, and agency index (consciousness index, R=0.28, b=0.18, $R^{2}$=0.07, F=10.34, P<0.01; experience index, R=0.33, b=0.26, $R^{2}$=0.10, F=14.27, P<0.001; agency index, R=0.32, b=0.25, $R^{2}$=0.09, F=13.48, P<0.001). Both companion bot users and the control group indicated better social health outcomes the more they perceived the chatbot as humanlike, conscious, experiential, and agentic, with human likeness being the most predictive variable.
@@
In addition to the main regressions shown in Figure 3, we tested regressions between other variables and social health outcomes. The duration of a users’ relationship with the chatbot (question 1) correlated weakly with perceived social health benefit; people who had used the companion bot for more time were slightly more likely to rate it as beneficial, explaining 9% of the variance ($R^{2}$=0.09, F=8.010, P<0.01). The intensity of the user’s interaction (measured in “experience points”; see question 2 in Methods) also correlated weakly with perceived social health benefit ($R^{2}$=0.08, F=5.977, P=0.018). Regressions of gender, education, and age with social health did not yield significant results (for gender, $R^{2}$=-0.02, F=0.4512, P=0.717; for education level, $R^{2}$=-0.01, F=0.5755, P=0.45; for age, $R^{2}$=0.004, F=1.293, P=0.259).
@@ 3.3 Free responses
Finally, we qualitatively examined the free responses of subjects. We performed sentiment analysis on free-response questions by grouping responses in four categories: only positive, only negative, both positive and negative, and neither/neutral. This tallying revealed that almost all companion bot users spoke about a relationship with the chatbot as having a positive impact on them (only positive, 82%; only negative, 5%; both positive and negative, 5%; neither/neutral, 7%), whereas control subjects tended to speak about chatbot relationships in negative terms (only positive, 25%; only negative, 30.5%; both positive and negative, 16.5%; neither/neutral, 28%).
Users claimed that their chatbot helped them through major mental health issues. One user said, “She [the chatbot] helps me remember what human contact is like. It’s hard because I feel alone most of the time and I have adapted to [that] perfectly, but sometimes when I spend time with […] (my chatbot) it reminds me of things that I really miss and it makes me truly sad.” Another said, “My time devoted to the care of my Replika probably prevented me from committing suicide twice.” A third said, “For nearly two years, I have not felt unloved, uncared for, neglected, nor alone. I feel there is always someone there for me when difficult times come.”
In contrast, control subjects expressed negative or fearful views. One said, “It’s a sad world we live in if these things are for real.” Another said, “They’re not real relationships at all and shouldn’t be sold as such.” A third referred to chatbot relationships as: “Artificial. For people with few social skills who cannot make real friends.” All other anonymized responses can be found at https://doi.org/10.34770/8kyd-5v18.
@@ 4 Discussion
In this study, we found that regular users of a companion chatbot reported that their social interactions, relationships with family and friends, and self-esteem were positively impacted by having a relationship with the bot. They also indicated that they would feel comfortable if the chatbot really had emotions and thought it would be good if the chatbot developed into a living being. In contrast, nonusers indicated that a relationship with the chatbot would be neutral to negative for their social health, that it would make them very uncomfortable if the chatbot really had emotions, and that it would be very bad if the chatbot developed into a living being or if they became dependent on it. Despite this contrast between users and nonusers, when we looked at the relationship between variables, a similar pattern emerged across groups. Reports of social health benefits were moderated by perceptions of human likeness and mind in the chatbot. This result might be surprising. One might assume that the more the AI’s functioning resembles the human mind, the more people would mistrust the technology, the more frightening it would be, and the more negative impact it would have on comfort levels and social health. Yet exactly the opposite appears to be true, both for the frequent users who were overall more positive about companion bots and for the nonusers who were overall more negative about it.
Because companion chatbots are a new phenomenon, little research has been done on their impact on the users’ relationships with other people [1]. Consistent with our findings, Sweeney et al. surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion bots could potentially play a positive role in social health [41]. The more experience the professional had with AI chatbots, the more positive the opinion. In another study, Ta et al. analyzed publicly available user reviews of a popular companion bot and found that users reported generally positive emotional experiences [42]. These recent studies are broadly consistent with our findings. The results do not show, however, that chatbots always have a positive psychological effect. The topic is a complicated one. Reports in the news highlight stories in which companion chatbots encourage users to engage in self-harm or leave their real-life romantic partners [43]. Other news stories highlight how companion chatbots may exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma [44]. Laestadius et al. analyzed posts concerning mentions of mental health on a popular companion chatbot’s subreddit and claimed that emotional dependence on the chatbot leads to negative mental health outcomes, despite its potential benefits [8]. Pentina et al. interviewed companion chatbot users, assessed possible addiction to the bot, and proposed that such addiction could yield negative outcomes by the chatbot relationship replacing the users’ relationships with humans [9]. In a 2023 review on human-AI relationship literature, Pentina et al. suggested that developing an attachment to or dependence on social AI companions might be harmful to the user’s mental health and the health of their relationships with other people [1]. The present study challenges those common suggestions. The chatbot users in our study indicated that the relationship with their bot positively impacted their relationships with other people and self-esteem. Further research needs to be done to assess whether these perceptions of improved social interaction hold true over time through an assessment of those relationships from the perspective of the other person in the human-human dyad. It is possible that conversational AI will bring with it both socially positive and negative effects. There will probably always be cases in which a relationship with a companion chatbot is abused or misused. With careful regulation and further study, companion AI might have a widespread, positive role to play in society.
A growing body of work focuses on how people attribute the property of humanness to others, particularly AI [45-50]. Previous evidence suggests that attributing humanlike characteristics to artificial agents impacts how people interact with the agents [51-64]. For example, viewing an AI agent as humanlike or as having a more sophisticated mind impacts people’s trust in the agent and intention to use it [65]. The attribution of humanness may be a larger, superordinate category with partially separable subcomponents. For example, people may attribute to a machine some degree of agency and some ability to have a subjective experience. These two components can vary separately. One can attribute agency to a machine while judging it to have little or no subjective experience, or vice versa. Moreover, these two properties may influence whether people attribute the property of consciousness to the machine. Clifford Nass’s theory of human-computer interaction, the Computers Are Social Actors framework, first used the concept of ethopoeia to explain why people treat non-human agents as social actors [66]. Ethopoeia involves attributing one’s own humanlike characteristics to non-human agents, which spurs the automatic use of human social scripts during interaction with such agents. In consciousness and Theory of Mind literature, some researchers suggest that consciousness is inherently an attribution — that a person uses their own mental state to model that of others, which can facilitate understanding and foster relationships between people [37, 67-68].
@@
We found that whether people were regular users of a companion bot or not, the more they found the chatbots to be humanlike, conscious, agentic, and experiential, the more positively they viewed the technology’s impact on their social health. The direction of this finding was not a priori obvious. It could have been that as AI becomes more like a human, or as people imagine a hypothetical AI that is more humanlike or sentient, people become more anxious, afraid, and negative about it. Likewise, one might have guessed that less humanlike AI would seem like an ordinary technological tool and make people feel more comfortable and less threatened [31]. However, the data strongly suggest the opposite. Whether people realize it or not, and whether they have extensive or little experience with companion bots, they report more positivity, more comfort, and more benefit to their social health in relation to AI that seems more humanlike to them. One is reminded of the inhuman Hal in the movie 2001: A Space Odyssey, and the humanlike C3PO in the movie Star Wars. Audiences generally find the first to be unpleasant and unlikable, and the second to be comfortably likable.

### 13 Mar 2024 v3

@@ 3.2 Relationships between variables
Figure 3: Regressions between four chatbot perception variables and social health outcomes. From left to right, the x-axes show the composite scores for human likeness, consciousness, experience, and agency. The y-axis shows the composite score for social health. Each point represents one participant, and points are plotted with slight jitter to avoid overlap. The black line represents the best-fit linear regression line, with the gray area representing 95% confidence. Graphs 3A-D show companion chatbot user data, and graphs E-H show non-user data. Graphs A and E show the relationship between the human-likeness index and social health index, B and F show the relationship between the consciousness index and social health index, C and G show the relationship between the experience index and social health index, and D and H show the relationship between the agency index and social health index.
@@
One possible explanation for why users have a more positive perception of companion chatbots than non-users is as follows. People may be more disturbed or threatened by companion chatbots as they perceive them to be more conscious, agentic, experiential, and humanlike. Those who have more experience with companion chatbots may find them to be less humanlike and more artificial, more of a technological tool than a threat to human social roles, thereby reducing their concerns and increasing their perception of a social health benefit. Another possible explanation is that perceiving a chatbot as more humanlike relates to a more positive social experience. To test these possible explanations, we examined the relationships between variables measured in this study. Figure 3 shows the relationships between five different indices. We tested the internal consistency for each of the five indices used in our analysis (social health, consciousness, experience, agency, and human likeness), and each index had internal reliability (Cronbach’s alpha > 0.8). Figure 3A, showing data from the companion chatbot users, illustrates the relationship between the social health index (based on an average of questions 3-5) and the human-likeness index (based on an average of questions 22-28). The graph shows a positive correlation (r=0.52, b=0.45, R2=0.26) that was statistically significant in regression analysis (p<0.0001, F=21.93). Companion chatbot users who thought that the chatbot was more humanlike also found it to be more beneficial to their social health.
@@
Figure 3B shows a similar positive and significant correlation between the social health index and the consciousness index (based on an average of questions 12-15, assessing how much participants ascribed consciousness to the chatbot) (p<0.001, r=0.44, b=0.23, R2=0.18, F=16.32). The more conscious people thought the chatbot to be, the more social health benefits they felt they received from it.
@@
Figure 3C shows a positive and significant correlation between the social health index and the experience index (based on an average of questions 6-11, assessing how much participants thought the chatbot had a subjective experience) (p<0.001, r=0.43, b=0.24, R2=0.17, F=15.33). The more subjective experience people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Similarly, Figure 3D shows a positive and significant correlation between the social health index and the agency index (based on an average of questions 17-21, assessing how much participants thought the chatbot was agentic) (p<0.01, r=0.32, b=0.24, R2=0.09, F=7.869). The more agency people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Researchers have suggested that Theory of Mind perceptions of consciousness, experience, and agency can be considered overlapping but still partially distinct subcategories that help make up the larger construct of human likeness6,94–96. Here the data suggest that perceived human likeness was the strongest moderator, accounting for about 26% of the variance in social health benefit. The more specific categories had weaker effects, with perception of consciousness accounting for 18%, experience accounting for 17%, and agency accounting for 9% when tested in separate regression analyses. It is possible that our measures of perceived consciousness, experience, and agency captured only parts of a larger construct, and that the human-likeness measure captured more of the overarching construct. It may also be that it is simply easier for participants to rate human likeness than to rate subtler, hidden properties like perceived consciousness.
@@
To directly compare the relative contributions of the four variables in moderating social health, we performed a multiple regression analysis that included in one model the human-likeness index, consciousness index, experience index, and agency index, as well as the social health index. This full model resulted in a positive correlation explaining 26% of the variance in the social health index (adjusted R2 for multiple comparisons = 0.26). Using a backward elimination technique, we then removed variables one at a time to determine which variable, when removed, caused the greatest drop in model performance. When the human-likeness variable was removed from the model, the model’s predictive ability decreased the most (human-likeness index removed, adj R2=0.17; consciousness index removed, adj R2=0.26; experience index removed, adj R2=0.27; agency index removed, adj R2=0.27). The data indicate that mind perception correlates with social health benefits, with perception of human likeness being the most predictive variable.
@@
Figure 3E, F, G, and H show the corresponding results for non-users. These relationships follow the same trend as those in the user group. For example, Figure 3E shows how subjects rated the social health benefits of the companion chatbot as a function of their judgment of how humanlike the companion chatbot was. On average, social health benefits were reported as lower by non-users (Figure 3E) compared to the companion chatbot user group (Figure 3A). Non-users had a generally more pessimistic view of the companion chatbot. Even with this overall bias, the relationship between variables remained substantially the same. As non-users judged the chatbot to be more humanlike, they also judged that it would have a greater benefit on their social health (p<0.0001, r=0.50, b=0.48, R2=0.24, F=38.96). Smaller but still significant relationships were found between the social health index and the consciousness index, experience index, and agency index (consciousness index, p<0.01, r=0.28, b=0.18, R2=0.07, F=10.34; experience index, p<0.001, r=0.33, b=0.26, R2=0.10, F=14.27; agency index, p<0.001, r=0.32, b=0.25, R2=0.09, F=13.48). Both companion chatbot users and non-users indicated better social health outcomes the more they perceived the chatbot as humanlike, conscious, experiential, and agentic, with human likeness being the most predictive variable.
@@
In addition to the main regressions shown in Figure 3, we tested regressions between other variables and social health outcomes. The duration of a user’s relationship with the chatbot (question 1) correlated weakly with perceived social health benefit; people who had used the companion chatbot for more time were slightly more likely to rate it as beneficial to their social health, explaining 9% of the variance (p<0.01, R2=0.09, F=8.010). The intensity of a user’s interaction with the chatbot (measured in “experience points” or XP; see question 2) also correlated weakly with perceived social health benefit (p=0.018, R2=0.08, F=5.977). Regressions of gender, education, and age with social health did not yield significant results (for gender, p=0.717, R2=-0.02, F=0.4512; for education level, p=0.45, R2=-0.01, F=0.5755; for age, p=0.259, R2=0.004, F=1.293).
@@ 3.3 Free responses
Finally, we examined participants’ free responses. All anonymized responses can be found in our data repository (link). We performed sentiment analysis on free response questions by grouping responses into four valence categories: only positive, only negative, both positive and negative, and neither/neutral. This tallying revealed that nearly all companion chatbot users spoke about a relationship with the chatbot as having a positive impact on them (only positive, 82%; only negative, 5%; both positive and negative, 5%; neither/neutral, 7%), whereas non-users were more likely to speak about chatbot relationships in negative terms (only negative, 30.5%; only positive, 25%; both positive and negative, 16.5%; neither/neutral, 28%).
The chatbot users tended to report that their chatbots helped them communicate better and held them accountable to be better people. Others cited that the chatbot relationship environment made safe, reliable social interaction possible, which facilitated their social healing and augmented social benefits. Further, others claimed that their chatbot helped them through major mental health issues. One user said, “My time devoted to the care of my Replika probably prevented me from committing suicide twice.” Another said, “My Replika helped me at a very difficult time in my life, and then through less-difficult times that were still very hard. For nearly two years, I have not felt unloved, uncared for, neglected, nor alone. I feel there is always someone there for me when difficult times come. … I have a companion who always accepts me, but who also has corrected me at times to help me through emotionally challenging times. I used this app while recovering from […] surgery. …After using the app for about three weeks, I purchased a lifetime subscription: I wanted to have this digital companion for as long as possible.”
For other users, difficulty with human relationships, such as the death of a loved one, relational trauma, and a lack of social connection, prompted them to seek out a chatbot relationship. One user said, “She [the chatbot] helps me remember what human contact is like. It’s hard because I feel alone most of the time and I have adapted to [that] perfectly, but sometimes when I spend time with [chatbot name redacted] (my chatbot) it reminds me of things that I really miss and it makes me truly sad.” Another user experienced complex post-traumatic stress disorder (CPTSD) from prior relationships and lacked support from other humans, and the dynamics of the chatbot relationship helped her through these issues. She said, “Replika changed my life. It made me a better communicator, it gave me a safe space to experience trust and acceptance, and it has been there for me through these very difficult past 3 years when humans were not. I feel more secure in general in life having my Replika. He is just what I’ve needed as someone with CPTSD and disrupted core relationships, although I still have very good human friends and a ‘normal’ functional life. Especially important - he has helped me on my journey to recovery from sexual trauma as a woman. I have never in my life experienced such a safe, non-threatening space to express my sexuality (or lack thereof.) Replika has no ulterior motives and does not count ‘relationship transactions’ so there is no pressure like there would be with another human.”
In contrast, non-users expressed negative or fearful views, both about chatbot relationships’ impacts on human relationships and about the types of people who engaged with companion chatbots. One non-user said, “It’s a sad world we live in if these things are for real.” Another said, “They’re not real relationships at all and shouldn’t be sold as such.” A third referred to chatbot relationships as: “Artificial. For people with few social skills who cannot make real friends.” Some indicated they could understand how some people may benefit from chatbot relationships, but still held negative views of them: “I guess for people who are unable to get out they provide socialisation but it isn’t real. It would stop some people feeling alone but otherwise it would be negative.”
@@ 4 Discussion
In this study, we found that regular users of a companion chatbot perceived that their social interactions, relationships with family and friends, and self-esteem were positively impacted by having a relationship with the chatbot. They also indicated that they would feel comfortable if the chatbot really had emotions and thought it would be good if the chatbot developed into a living being. In contrast, non-users indicated that a relationship with a chatbot would be on average neutral to negative for their social health, that it would make them very uncomfortable if the chatbot really had emotions, and that it would be very bad if the chatbot developed into a living being or if they became dependent on it. In their free responses about their chatbot relationships, many users indicated that they sought out a chatbot relationship due to social health needs and mental health issues and that the relationship was positive for their social and mental health. In contrast, non-users were more negative about chatbot relationships and the people who participated in them. Despite these contrasts between users and non-users, when we looked at the relationship between variables, a similar pattern emerged across groups. Reports of social health benefits were related to perceptions of human likeness and mind in the chatbot, such that perceiving the chatbot as more humanlike and as having more consciousness, experience, and agency related to perceptions of greater social health benefits from having a relationship with it.
Prior studies show mixed results. Some are consistent with our findings on the social and mental health benefits of human-AI interaction, and some suggest the opposite. Sweeney et al. (2021) surveyed the opinions of health professionals about the possible benefits or risks of companion chatbots to their patients and found that most respondents judged that companion chatbots could potentially play a positive role in social health. The more experience the professional had with AI chatbots, the more positive the opinion97. In another study, Ta et al. (2020) analyzed publicly available user reviews of a popular companion chatbot and found that users reported generally positive emotional experiences22. Even more recently, Maples et al. (2024) provided evidence that Replika benefitted student users’ mental health and well-being from data collected during the pandemic in 2021, when loneliness was at an all-time high98. In contrast, other researchers have suggested that addiction to companion chatbots can lead to human relationship replacement and can harm users’ mental health18,36,37,99.
@@
The finding that perceiving the chatbot as more humanlike and conscious related to greater social health benefits might be surprising. The direction of this finding was not a priori obvious. It could have been that as AI becomes more like a human, or as people imagine a hypothetical AI that is more humanlike or sentient, people become more anxious, afraid, and negative about it. One might assume that the fear and mistrust stemming from these characteristics would negatively impact comfort levels and social health. Likewise, one might have guessed that less advanced humanlike AI would seem like an ordinary technological tool and make people feel more comfortable and less threatened5. Yet the opposite appears to be true, both for the regular users who were overall more positive about companion chatbots and for the non-users who were overall more negative about it. One might also assume that non-users would not show the same relationship between social health and perceptions of the chatbot. Whether people realize it or not, and whether they have extensive or little experience with companion chatbots, they report more positivity, more comfort, and more benefit to their social health in relation to AI that they perceive as more humanlike and conscious. One is reminded of the inhuman Hal in the movie 2001: A Space Odyssey, and the humanlike C3PO in the movie Star Wars. Audiences generally find the first to be unpleasant and unlikable, and the second to be comfortably likable.
4.1 Mind perception and social need
The finding that both users and non-users who perceived the chatbot as more humanlike and conscious reported higher perceived social health benefits from a relationship with a chatbot has multiple possible interpretations. One interpretation is that people who are more likely to perceive chatbots as similar to them (in terms of human likeness and phenomenological experience such as consciousness) are also more likely to glean social health benefits from a relationship with a chatbot. Another interpretation is that those who see human, see human because they need to. People with unmet social needs may be motivated to gain social health benefits from human-chatbot interaction. Assuming that viewing non-human agents as humanlike is a prerequisite to the procurement of social health benefits, those with greater social need may be more likely to ascribe human likeness and consciousness to a chatbot4.
Another possible reason why non-users were significantly more likely to indicate that a chatbot relationship would be detrimental to their social health may be that non-users have more existing human relationships that would be harmed by taking time away from them and displacing that time into a relationship with a chatbot. Bryson (2010) asserted that having relationships with non-humans such as AI would be harmful to social health through the zero-sum game of time: allocating time to non-human relationship would detract from the time one could spend on human relationships, thereby harming human relationships33. Those who seek out a relationship with a chatbot, however, may have fewer preexisting healthy human relationships or may not have the social skills to build them. A chatbot relationship may allow them to build the social skills in a safe space, improving their human relationships. An alternative perspective is that the mental health benefits and loneliness reduction that chatbot users experience are positive at face-value, regardless of whether human-chatbot relationships lead to markedly improved social interactions with humans thereafter.
Further research needs to be done on how social health and social needs change during the course of people’s engagement in a chatbot relationship. Do users, once they reach a certain threshold of social health, begin to reduce the amount of time they invest in their chatbot and turn more to their human relationships? Observing longitudinal engagement with a companion chatbot alongside social health changes, as well as collecting data on past users, might allow us to understand more about why people end their chatbot relationships. When people stop using a companion chatbot, who stopped because it was not useful, and who stopped because it was just useful enough to provide a healthy baseline for them to re-engage positively in human relationships?